You’re working at a biochemistry lab. A professor

asked you to prepare 1.0 L of 0.10 M acetate buffer

of pH 4.00. She told that you can prepare it by

mixing 0.10 M solution of acetic acid (pKa 4.76) and

0.10M of sodium acetate solution. How much

volume of 0.10 M solutions of acetic acid (X) and

sodium acetate (Y) do you need respectively?
Answer #1

Solution:

According to Hendersen-Hasselbalch equation,

pH = pKa + log [CH3COONa) ] / [CH3COOH]

4.00 = 4.76 + log [CH3COONa]/[CH3COOH]

-0.76 = log [CH3COONa] / [CH3COOH]

[CH3COONa]/ [CH3COOH) = antilog (-0.76) = 0.174

Thus,

[CH3COONa] = 0.174 x [CH3COOH]

For each volume of acetic acid, 0.174 volume of sodium acetate is added, therefore,

1 + 0.174 = 1.174

Thus,

Volume of acetic acid =( 1/1.174) x 1000 mL = 851.8 mL

Volume of acetate = (0.174 /1.174) x 1000 mL = 148.2 mL

Hence, X = 851.8 mL

Y = 148.2 mL